the science department purchases sulphuric acid in 1.0L bottles with a concentration of 18.0M. How much acid (mL) is needed to make 100.0mL of 15.0M acid?
1
Expert's answer
2020-04-23T13:09:13-0400

n(HCl) = 100/1000*15 = 1.5 mol - in final solution, so amount of HCl taken from initial soultion must be the same.

V1 = volume of initial solution

V1*18 = 1.5 mol

V1 = 1.5/18 =0.08333 L = 83.33 mL